Use the behavioural equations below and answer the questions that follow:

C = R4 billion + c1YD = R19.7 billion

Y= R13.1 billion

T= R6 billion

c0 = R13.5 billion


i.                   Calculate the propensity to consume.                                                                 (4)



1
Expert's answer
2021-06-29T08:15:36-0400

Here we are given the consumption equation C = 4billion + c1YD = 19.7billion

Y means income so here we have Y = 13.1billion

T refers to taxes and here T = 6billion

c0 is autonomous consumption and here it is c0 = 13.5billion

Solution is as follows :- 

          "C = c_0 + cY_d"

 "19.7=13.5 + c (Y \u2212 T)"  since Yd is disposable income ( income − taxes )      

         

"19.7=13.5 + c (13.1 − 6) \\ 

 19.7=13.5 + c (7.1) \\

 19.7-13.5=c (7.1) \\ 

 6.2=c 7.1 \\

 c=\frac{6.2 }{ 7.1}\\ 

 c=0.87"

Therefore we have the marginal propensity to consume as 0.87